# Intramolecular Singlet Fission Through a Coherently Coupled Excimer-like Intermediate

### v1

@@ 2.3 $TT_{1}$ formation through a coherently coupled intermediate
To further understand the 200 fs rise and decay of the ESA signal at the detection wavelengths of 710 nm and 760 nm respectively, we ignore the above excitation wavelength dependence and globally fit the 2D spectra with a freely floated three-exponential model. The corresponding rate maps are shown in Figure 2c. A dispersive lineshape along the detection wavelength region is seen for the $\tau_{1}$=0.19 $\pm$ 0.02 ps signal which confirms the visual trends seen in Figure 2b. As may be expected from the neglect of wavelength-dependent dynamics in the global rate map analysis, the $\tau_{2}$=1.80 $\pm$ 0.21 ps rate map shows a clear trend in the 650 nm excitation wavelength region ( Figure S3c). Interestingly, hardly any rise is seen in the positive GSB band which is consistent with the observation of positive CPU between the upper and lower Davydov components ( Figure 2a). This is so because the GSB signal strength37, 38, 39, $s^{GSB}$, in a strongly electronically coupled system is already twice that of stimulated emission (SE), $s^{SE}$, because shared electronic correlations (shared ground state) implies that bleaching one transition also bleaches the other. Therefore, in this system, no rise in the GSB signal strength during $LE\rightarrow TT_{1}$ conversion is to be expected.
The signal-to-noise ratio (SNR) of the 2D signal is $\sim$2.42 due to low pump pulse energies ($\sim$2.1 nJ over 200 nm) and sample extinction coefficient ($\varepsilon$ $\sim$5530 M-1cm-1). Therefore we confirm the above trends from a separately conducted impulsive PP experiment with better SNR. This also allows us to track the impulsively excited quantum beats, if any, as reporters40, 41 for the rich excited state dynamics observed from 2DES measurements. The PP measurements and the corresponding target analysis are presented in Figure 3.
@@
Figure 3a shows the PP spectra as a function of waiting time $T$. The 200 fs rise (fall) of the signal at blue (red) detection wavelengths, marked by dashed versus solid yellow squares in Figure 2b, are evident from the blue and red ESA bands in Figure 3a. The corresponding band-integrated $T$ dynamics is shown in Figure 3b, which shows the concomitant rise/fall dynamics. A global fit of the pump-probe data reveals 0.19 $\pm$ 0.02 ps and 1.80 $\pm$ 0.21 ps as the fastest two time constants, both in good agreement with the 2DES data if the excitation wavelength dependence is ignored. Expectedly, when this global time constant is imposed on the 2D rate maps, a clear trend in the residual is seen for excitation wavelengths corresponding to the lower Davydov component around 650 nm. This is shown in Figure S3b,c. The longest time constant of 73.58 $\pm$ 6.58 ps is in good agreement with the 70 ps time constant reported earlier28 and assigned to triplet-triplet geminate recombination. The details of the global fit are summarized in Section S1.2.2. Figure 3c,d show that a minimal target model which allows for the initially excited singlet $S_{1}$ to be converted to an intermediate which then relaxes to the $TT_{1}$ state satisfactorily fits the PP data. The concentration profile in Figure 3d shows that the population of the intermediate species is already maximized by 480 fs. The corresponding species associated spectra (SAS) in Figure 3c show only minimal spectral changes between the initially excited $S_{1}$ and the $TT_{1}$ photoproduct species with strongly overlapping ESA bands. Note that in the earlier pump-probe study28 the $TT_{1}$ species was confirmed through a Ruthenium-based triplet sensitizer that transferred triplets to the NDI dimer whose triplet spectrum was measured through a microsecond pump-probe experiment and found to be quite similar to that of $TT_{1}$. The similarity of the spectrum of the intermediate species to the initial $S_{1}$ and the final $TT_{1}$ photoproduct suggests a coherent intermediate formation, as illustrated in Figure 3e, with minimal electronic reorientation during the nuclear evolution away from the initial FC geometry. This assertion is also consistent with the fast formation time of the intermediate species, its excitation wavelength-independent formation rate ( Figure 2b), and minimal electronic reorientation during this process (vide infra). We will come back to this latter point during the discussion of electronic polarization anisotropy.
2.4 Tracking electronic reorientation during $S_{1}\rightarrow[I]\rightarrow TT_{1}$
@@
A priori, during internal conversion between species with orthogonally polarized transition dipoles, such as $S_{1}\rightarrow TT_{1}$ conversion in acenes42 with transition dipole $\hat{\mu}_{S_{0}\rightarrow S_{1}}$ along the short axis and $\hat{\mu}_{TT_{1}\rightarrow TT_{n}}$ polarized along the long axis, anisotropy43 $r(\theta_{ij})=(3\cos^{2}(\theta_{ij})-1)/{5}$, where $i,j$ are the initial and final species and $\theta$ is the relative angle between their transition dipoles, can be expected to change from 0.4 to -0.2. A similar anisotropy of approximately -0.2 has been recently reported44 by Vauthey and co-workers in the context of a $CT$ donor-acceptor system. In the current context, species $i$ is the initially excited singlet while $j$ could be the intermediate or later the relaxed $TT_{1}$ species. The above changes in anisotropy can be directly related to $\Delta\theta_{ij}$ change of $3.53^{\circ}$ and $3.49^{\circ}$ at 600 fs and 6 ps, respectively, in the electronic transition dipole orientation relative to 200 fs, suggesting minimal electronic reorientation on the excited state during iSF. However, a minimal change in the anisotropy in the GSB/SE band may not reflect the true electronic reorientation because of weak SE signal strength. This is so because ultrafast quenching of the SE signal due to iSF implies that the total anisotropy $r_{tot}$ in the GSB/SE band, given by $r_{tot}=(s^{GSB}\cdot r_{GSB}+s^{SE}\cdot r_{SE})/({s^{GSB}+s^{SE}})$, will be sensitive to the GSB signal only. In contrast, anisotropy in the ESA band, shown in Figure 4b (right), reflects excited state changes and provides the opportunity to exclusively track electronic reorientation during $S_{1}\rightarrow TT_{1}$ conversion without any contamination from the ground state anisotropy. Surprisingly, minimal changes in the anisotropy are observed in the ESA band as well. The expected anisotropy change during $S_{1}\rightarrow TT_{1}$ conversion, that is parallel to orthogonally polarized initial and final states transitions, $\mu_{S_{0}\rightarrow S_{1}}\parallel\mu_{S_{1}\rightarrow S_{n}}$ and $\mu_{S_{0}\rightarrow S_{1}}\perp\mu_{TT_{1}\rightarrow TT_{n}}$, respectively, is 0.4 to -0.2. Compared to this, the average $r_{ESA}$ in the three time windows is $0.145\pm 0.013$, $0.150\pm 0.014$ and $0.152\pm 0.016$, respectively ( Table S3), all within the error bars of each other. Note that we have confirmed that the anisotropy in the ESA band stays $\sim$0.15 with $T$ across the band and not sensitive to the position of the chosen band. This observation directly reports negligible electronic reorientation during the nuclear evolution towards the intermediate $[I]$ and its relaxation to the $TT_{1}$ state. It also raises the question that what causes the observed ESA anisotropy to deviate so much from the expected limits of the anisotropy during iSF ? A similar question also arises for the GSB/SE band anisotropy of 0.215 at 200 fs which lies significantly below the expected 0.4 for an isolated transition dipole.
@@
To better understand the limits of polarization anisotropy observed in the experiments, results37 from purely excitonic dimers with orthogonal transition dipoles provide a useful starting point. Although the case of iSF dimer is more complicated due to $CT$ and $TT_{1}$ states which may borrow oscillator strength from the $LE$ states, in the impulsive limit where all states are excited, $r_{GSB}$ of both dimers will be the same because oscillator strength arises from the $LE$ states in both cases. Strongly correlated electronic states with orthogonal transition dipoles which share a common ground state, such that the signal strengths are $s^{GSB}=2s^{SE}$, show37, 45 $r_{GSB}$ between 0.1 in the impulsive excitation limit to 0.4 in the single transition dipole excitation limit. In comparison, the SE signal, assuming no electronic coherence signal pathways, shows $r_{SE}$ ($T=0$) of 0.4 corresponding to SE from a transition dipole that is maximally aligned with the pump polarization. Section S4 extends these results to a generalized excitonic dimer with any angle $\theta_{AB}$ between the transition dipoles associated with sites $A$ and $B$.
@@
In case of orthogonal dipoles, the GSB anisotropy monotonically increases45 from 0.1 under impulsive excitation to 0.4 under (non-impulsive) single transition dipole excitation. For the generalized dimer model, Figure S5 shows that the monotonic increase45 in $r_{GSB}$ from 0.1 to 0.4 is reversed at intermediate values of $\theta_{AB}$. This is due to the interference between the transition dipoles which is absent for the orthogonal case. For $\theta_{AB}$ = 24o, the expected GSB anisotropy with impulsive excitation that covers all states is $r_{GSB}$ = 0.35 compared to the observed value of 0.215 at 200 fs. From Figure S5, this can be understood as due to the partial laser coverage of the upper Davydov component at $\sim$600 nm as well as to its loss of oscillator strength that gets distributed to higher lying $CT$ states due to large LECTLE-CT mixing. As we noted earlier, indeed the TD-DFT calculations36 from Zhu and co-workers on a contorted PDI dimer show that the singlet oscillator strength is significantly re-distributed among higher-lying $CT$ and singlet configurations (see Section S8 of ref. 36). The observation of low GSB anisotropy could also mean that the relative orientation of the NDI chromophores in the dimer changes significantly between the crystal and the solution. This possibility seems unlikely for a rigid dimer. Note that even if the SE signal was considered in the estimation of GSB/SE band anisotropy, the total anisotropy will only increase to deviate more from the observed anisotropy. Figure S7 estimates the amount of ESA contribution in the GSB/SE band which shows only $\sim 10\%$ expected contribution thus also not explaining the reduced ansisotropy in the GSB/SE band. Overall, reduced anisotropy in the GSB/SE band complements the evidence of strongly coupled iSF dimer with shared ground state correlations reported by the positive upper cross peak from 2DES ( Figure 2)
@@
As we noted earlier, the ESA band provides the opportunity to look at purely excited state dynamics. The expected range of $r_{ESA}$ from 0.4 to -0.2, for 0% to 100% $\hat{\mu}_{TT_{1}\rightarrow TT_{n}}$ character (100% to 0% $\hat{\mu}_{S_{1}\rightarrow S_{n}}$ character) of the ESA band, respectively, suggests that the experimentally observed $r_{{ESA}}$ of $\sim 0.15$ represents an interesting case where the ESA band is polarized in between $\hat{\mu}_{S_{1}\rightarrow S_{n}}$ and $\hat{\mu}_{TT_{1}\rightarrow TT_{n}}$ transitions. This is not surprising given the nearly identical spectra of the singlet and the triplet species (Figure3). Assuming both transitions are present within the ESA band with strengths $\kappa$ and $\gamma$, respectively, the total anisotropy in the ESA band can be expressed46 as $r_{ESA}=(\kappa^{2}r_{S_{1}\rightarrow S_{n}}+\gamma^{2}r_{TT_{1}\rightarrow TT_{n}})/({\kappa^{2}+\gamma^{2}})$. Equating this to the observed anisotropy predicts that the ESA band consists of $\sim$60-40 ratio of $\hat{\mu}_{S_{1}\rightarrow S_{n}}$ and $\hat{\mu}_{TT_{1}\rightarrow TT_{n}}$ polarized transitions. The intermediate polarization of the ESA band with no significant change with $T$ directly reports that the $S_{1}\rightarrow[I]\rightarrow TT_{1}$ internal conversion process involves surprisingly minimal electronic reorientation with strong electronic correlations between the chromophore sites preserved during the internal conversion process.
[…]
@@
Figure 5: Enhanced quantum beats in the $TT_{1}$ photoproduct correspond to inter-chromophore nuclear motions. (a) Quantum beat spectrum as a function of detection wavelength. The top panel shows the PP spectrum at $T$ = 1 ps, while the side panel shows the ground state Raman spectrum of the NDI dimer excited at 785 nm along with the non-resonant Raman spectra of the two solvents ACN and DCM. The lowest contour is at 20%. Quantum beats in the PP data arising from the non-resonant Raman scattering from the solvent are marked as the horizontal grey band. (b) The vertical gray bands in panel a are spectrally integrated and overlaid for the GSB/SE (blue) and ESA (red) bands. Vertical gray bands mark the solvent bands, same as horizontal bands in panel a. (c) Schematic for the excimer-like intermediate formation due to large excited state nuclear displacements along the theoretical 219 cm-1 and 590 cm-1 modes that correspond to inter-chromophore ruffling and twisting motions, respectively.
@@
Figure 5a shows the quantum beat spectrum as a function of detection wavelength obtained from the PP data in Figure 3. The PP spectra shows two modes at 210 cm-1 and 575 cm-1 that are prominently seen on the $TT_{1}$ photoproduct and nearly absent in the ground electronic state. This can also be seen in Figure 5b which overlays the spectrally integrated GSB and ESA bands which are denoted by the gray vertical bands in Figure 5a. The ground state Raman spectrum for the NDI dimer, overlaid as left vertical panel, shows only weak Raman peaks corresponding to the quantum beats at $\sim$210 cm-1 and $\sim$575 cm-1 and consistent with their near absence in the GSB band observed in the PP experiments. Further, Figure S4a shows a similar plot as Figure 5a for the NDI monomer. Exactly similar trends are seen in the NDI monomer as well, that is, very weak to absent modes at $\sim$180 cm-1 and $\sim$580 cm-1 on the ground electronic state, confirmed by Raman spectra as well as the PP spectra. The absence of these prominent modes on the ground electronic state of the NDI monomer is also consistent with the Raman spectrum of related NDI derivatives reported47 previously (see Figure 5 of ref. 47). Figure S4c also compares the mode frequencies between the NDI monomer and dimer where shifts of 30 cm-1 for the 210 cm-1 mode and 5 cm-1 for the 575 cm-1 are seen. These shifts likely arise due to the constrained skeleton of the dimer.

### v3

@@ 2.3 $TT_{1}$ formation through a coherently coupled intermediate
To further understand the 200 fs rise and decay of the ESA signal at the detection wavelengths of 710 nm and 760 nm respectively, we ignore the above excitation wavelength dependence and globally fit the 2D spectra with a freely floated three-exponential model. The corresponding rate maps are shown in Figure 2c. A dispersive lineshape along the detection wavelength region is seen for the $\tau_{1}$=0.19 $\pm$ 0.02 ps signal which confirms the visual trends seen in Figure 2b. As may be expected from the neglect of wavelength-dependent dynamics in the global rate map analysis, the $\tau_{2}$=1.80 $\pm$ 0.21 ps rate map shows a clear trend in the 650 nm excitation wavelength region ( Figure S3c). Curiously, hardly any rise is seen in the positive GSB band, which is indeed consistent with the observation of positive CPU between the upper and lower Davydov components ( Figure 2a). This is so because the GSB signal strength41, 42, 43, $s^{GSB}$, in a strongly electronically coupled system is already twice that of stimulated emission (SE), $s^{SE}$ – shared electronic correlations (shared ground state) implies that bleaching one transition also bleaches the other. Therefore, for a system with positive, upper 2D cross-peak, no rise in the GSB signal during $LE\rightarrow TT_{1}$ conversion is to be expected.
The signal-to-noise ratio (SNR) of the 2D signal is $\sim$2.42 due to low pump pulse energies ($\sim$2.1 nJ over 200 nm) and sample extinction coefficient ($\varepsilon$ $\sim$5530 M-1cm-1). Therefore we confirm the above trends from a separately conducted impulsive PP experiment with better SNR. This also allows us to track the impulsively excited quantum beats, if any, as reporters44, 45 for the rich excited state dynamics observed from 2DES measurements. The PP measurements and the corresponding target analysis are presented in Figure 3.
@@
Figure 3a shows the PP spectra as a function of waiting time $T$. The 200 fs rise (fall) of the signal at blue (red) detection wavelengths, marked by dashed versus solid yellow squares in Figure 2b, are evident from the blue and red ESA bands in Figure 3a. The corresponding band-integrated $T$ dynamics is shown in Figure 3b, which shows the concomitant rise/fall dynamics. A global fit of the pump-probe data reveals 0.19 $\pm$ 0.02 ps and 1.80 $\pm$ 0.21 ps as the fastest two time constants, both in good agreement with the 2DES data if the excitation wavelength dependence is ignored. Expectedly, when this global time constant is imposed on the 2D rate maps, a clear trend in the residual is seen for excitation wavelengths corresponding to the lower Davydov component around 650 nm. This is shown in Figure S3b,c. The longest time constant of 73.58 $\pm$ 6.58 ps is in good agreement with the 70 ps time constant reported earlier32 and assigned to triplet-triplet geminate recombination. The details of the global fit are summarized in Section S1.2.2. Figure 3c,d show that a minimal target model which allows for the initially excited singlet $S_{1}$ to be converted to an intermediate which then relaxes to the $TT_{1}$ state satisfactorily fits the PP data. The concentration profile in Figure 3d shows that the population of the intermediate species is already maximized by 480 fs. The corresponding species associated spectra (SAS) in Figure 3c show only minimal spectral changes between the initially excited $S_{1}$ and the $TT_{1}$ photoproduct species with strongly overlapping ESA bands. Note that in the earlier pump-probe study32 the $TT_{1}$ species was confirmed through a Ruthenium-based triplet sensitizer that transferred triplets to the NDI dimer whose triplet spectrum was measured through a microsecond pump-probe experiment and found to be quite similar to that of $TT_{1}$. The similarity of the spectrum of the intermediate species to the initial $S_{1}$ and the final $TT_{1}$ photoproduct has been reported46, 27 previously in case of intermolecular SF. Combined with our other observations, this suggests that the coherent $S_{1}+TT_{1}]$ intermediate, denoted as $[I]$ in the Figure 3, forms with minimal electronic reorientation during the nuclear evolution away from the initially excited FC geometry. This assertion is also consistent with the fast formation time of the intermediate species, its excitation wavelength and solvent polarity independent formation rate ( Figure 2b and Section S5, respectively), and minimal electronic reorientation during this process (vide infra). We will come back to this latter point during the discussion of electronic polarization anisotropy. The overall picture of coherent intermediate formation emerging from the above discussion is summarized in Figure 3e.
2.4 Tracking electronic reorientation during $S_{1}\rightarrow[S_{1}+TT_{1}]\rightarrow TT_{1}$
@@
A priori, during internal conversion between species with orthogonally polarized transition dipoles, such as $S_{1}\rightarrow TT_{1}$ conversion in acenes47 with transition dipole $\hat{\mu}_{S_{0}\rightarrow S_{1}}$ along the short axis and $\hat{\mu}_{TT_{1}\rightarrow TT_{n}}$ polarized along the long axis, anisotropy48 $r(\theta_{ij})=(3\cos^{2}(\theta_{ij})-1)/{5}$, where $i,j$ are the initial and final species and $\theta$ is the relative angle between their transition dipoles, can be expected to change from 0.4 to -0.2. A similar anisotropy of approximately -0.2 has been recently reported49 by Vauthey and co-workers in the context of a $CT$ donor-acceptor system. In the current context, species $i$ is the initially excited singlet while $j$ could be the intermediate or later the relaxed $TT_{1}$ species. The above changes in anisotropy can be directly related to $\Delta\theta_{ij}$ change of $3.53^{\circ}$ and $3.49^{\circ}$ at 600 fs and 6 ps, respectively, in the electronic transition dipole orientation relative to 200 fs, suggesting minimal electronic reorientation on the excited state during iSF. However, a minimal change in the anisotropy in the GSB/SE band may not reflect the true electronic reorientation because of weak SE signal strength. This is so because ultrafast quenching of the SE signal due to iSF implies that the total anisotropy $r_{tot}$ in the GSB/SE band, given by $r_{tot}=(s^{GSB}\cdot r_{GSB}+s^{SE}\cdot r_{SE})/({s^{GSB}+s^{SE}})$, will be sensitive to the GSB signal only. In contrast, anisotropy in the ESA band, shown in Figure 4b (right), reflects excited state changes and provides the opportunity to exclusively track electronic reorientation during $S_{1}\rightarrow TT_{1}$ conversion without any contamination from the ground state anisotropy. Surprisingly, minimal changes in the anisotropy are observed in the ESA band as well. The expected anisotropy change during $S_{1}\rightarrow TT_{1}$ conversion, that is parallel to orthogonally polarized initial and final states transitions, $\mu_{S_{0}\rightarrow S_{1}}\parallel\mu_{S_{1}\rightarrow S_{n}}$ and $\mu_{S_{0}\rightarrow S_{1}}\perp\mu_{TT_{1}\rightarrow TT_{n}}$, respectively, is 0.4 to -0.2. Compared to this, the average $r_{ESA}$ in the three time windows is $0.145\pm 0.013$, $0.150\pm 0.014$ and $0.152\pm 0.016$, respectively ( Table S3), all within the error bars of each other. Note that we have confirmed that the anisotropy in the ESA band stays $\sim$0.15 with $T$ across the band and not sensitive to the position of the chosen band. This observation directly reports negligible electronic reorientation during the nuclear evolution towards the intermediate $S_{1}+TT_{1}]$ and its subsequent relaxation to the $TT_{1}$ state. It also raises the question that what causes the observed ESA anisotropy to deviate so much from the expected limits of the anisotropy during iSF ? A similar question also arises for the GSB/SE band anisotropy of 0.215 at 200 fs which lies significantly below the expected 0.4 for an isolated transition dipole.
@@
To better understand the limits of polarization anisotropy observed in the experiments, results41 from purely excitonic dimers with orthogonal transition dipoles provide a useful starting point. Although the case of iSF dimer is more complicated due to $CT$ and $TT_{1}$ states which may borrow oscillator strength from the $LE$ states, in the impulsive limit where all states are excited, $r_{GSB}$ of both dimers will be the same because oscillator strength arises from the $LE$ states in both cases. Strongly correlated electronic states with orthogonal transition dipoles which share a common ground state, such that the signal strengths are $s^{GSB}=2s^{SE}$, show41, 50 $r_{GSB}$ between 0.1 in the impulsive excitation limit to 0.4 in the single transition dipole excitation limit. In comparison, the SE signal, assuming no electronic coherence signal pathways, shows $r_{SE}$ ($T=0$) of 0.4 corresponding to SE from a transition dipole that is maximally aligned with the pump polarization. Section S4 extends these results to a generalized excitonic dimer with any angle $\theta_{AB}$ between the transition dipoles associated with sites $A$ and $B$.
@@
In case of orthogonal dipoles, the GSB anisotropy monotonically increases50 from 0.1 under impulsive excitation to 0.4 under (non-impulsive) single transition dipole excitation. However, for the generalized dimer model, Figure S5 shows that the monotonic increase50 in $r_{GSB}$ from 0.1 to 0.4 is reversed at intermediate values of $\theta_{AB}$. This is due to the interference between the transition dipoles which is absent for the orthogonal case. For $\theta_{AB}$ = 24o, the expected GSB anisotropy with impulsive excitation that covers all states is $r_{GSB}$ = 0.35 compared to the observed value of 0.215 at 200 fs. From Figure S5 the reduced anisotropy can be understood as due to the partial laser coverage of the upper Davydov component at $\sim$600 nm as well as to its loss of oscillator strength that gets distributed to higher lying $CT$ states due to large LECTLE-CT mixing. As we noted earlier, indeed the TD-DFT calculations36 from Zhu and co-workers on a contorted PDI dimer show that the singlet oscillator strength is significantly re-distributed among higher-lying $CT$ and singlet configurations (see Section S8 of ref. 40). The observation of reduced GSB anisotropy could also mean that the relative orientation of the NDI chromophores in the dimer changes significantly between the crystal and the solution. This possibility seems unlikely for a rigid dimer. Note that even if the SE signal was considered in the estimation of GSB/SE band anisotropy, the total anisotropy will only increase to deviate more from the observed anisotropy. Figure S7 shows that the amount of ESA contribution in the GSB/SE band is only $\sim 10\%$ and not sufficient to explain the reduced ansisotropy in the GSB/SE band. Overall, reduced anisotropy in the GSB/SE band complements the evidence of strongly coupled iSF dimer with shared ground state correlations, consistent with the positive, upper cross peak reported by the 2DES spectra ( Figure 2).
@@
As we noted earlier, the ESA band provides the opportunity to look at purely excited state dynamics. The expected range of $r_{ESA}$ from 0.4 to -0.2, for 0% to 100% $\hat{\mu}_{TT_{1}\rightarrow TT_{n}}$ character (100% to 0% $\hat{\mu}_{S_{1}\rightarrow S_{n}}$ character) of the ESA band, respectively, suggests that the experimentally observed $r_{{ESA}}$ of $\sim 0.15$ represents an interesting case where the ESA band is polarized in between $\hat{\mu}_{S_{1}\rightarrow S_{n}}$ and $\hat{\mu}_{TT_{1}\rightarrow TT_{n}}$ transitions. This is not surprising given the nearly identical spectra of the singlet and the triplet species (Figure3). Assuming both transitions are present within the ESA band with strengths $\kappa$ and $\gamma$, respectively, the total anisotropy in the ESA band can be expressed51 as $r_{ESA}=(\kappa^{2}r_{S_{1}\rightarrow S_{n}}+\gamma^{2}r_{TT_{1}\rightarrow TT_{n}})/({\kappa^{2}+\gamma^{2}})$. Equating this to the observed anisotropy predicts that the ESA band consists of $\sim$60-40 ratio of $\hat{\mu}_{S_{1}\rightarrow S_{n}}$ and $\hat{\mu}_{TT_{1}\rightarrow TT_{n}}$ polarized transitions. The intermediate polarization of the ESA band with no significant change with $T$ directly reports that the $S_{1}\rightarrow[S_{1}+TT_{1}]\rightarrow TT_{1}$ internal conversion process involves surprisingly minimal electronic reorientation, maintains the strong electronic correlations between the chromophore sites with significant singlet-triplet mixing throughout $TT^{1}$ formation and relaxation. To best of our knowledge, the surprising lack of electronic reorientation during iSF has not been reported before and suggests that synthetic design should aim for minimizing such mixing to prolong the lifetime of high-spin correlated triplets. Such mixing also needs to be accounted for in the electronic structure methods that aim to predict the high-spin state dynamics that follows $(TT_{1})$ generation.
At this point, the question arises as to what is the nature of the intermediate $S_{1}+TT_{1}]$. Measurements of weakly emissive sharp emission profiles in the excitation–emission spectra ( Figure 1b,c) suggest intensity borrowing arising from LECTLE-CT mixing is likely. However, the position of these weakly emissive states is at least 2043 cm-1 (75 nm) above the lower Davydov component at 15625 cm-1 (640 nm). Signatures of NDI radical cation and anion bands32, 33 in the PP spectra are absent. Figure S8 overlays the cation and anion bands with the NDI dimer PP spectrum where no such bands are obvious. Thus no significant $CT$ character may be expected in the intermediate state. This assertion is further supported by our observations of solvent polarity independent rates ( Figure S6) for both steps – the intermediate formation as well as its wavelength dependent relaxation to $TT_{1}$. SF through intermediates with minimal to absent $CT$ character have been reported46, 26 previously. To further investigate this question, we measured impulsive vibrational quantum beat spectra during the $S_{1}$$S_{1}+TT_{1}]$$TT_{1}$ electronic relaxation.
@@
Figure 5: Enhanced quantum beats in the $TT_{1}$ photoproduct correspond to inter-chromophore nuclear motions. (a) Quantum beat spectrum as a function of detection wavelength. The top panel shows the PP spectrum at $T$ = 1 ps, while the side panel shows the ground state Raman spectrum of the NDI dimer excited at 785 nm along with the non-resonant Raman spectra of the two solvents ACN and DCM. The lowest contour is at 20%. Quantum beats in the PP data arising from the non-resonant Raman scattering from the solvent are marked as the horizontal grey band. (b) The vertical gray bands in panel a are spectrally integrated and overlaid for the GSB/SE (blue) and ESA (red) bands. Vertical gray bands mark the solvent bands, same as horizontal bands in panel a. (c) Schematic for the excimer-like intermediate formation due to large excited state nuclear displacements along the theoretical 219 cm-1 and 590 cm-1 modes that correspond to inter-chromophore ruffling and twisting motions, respectively (Section S1,S3).
@@
Figure 5a shows the quantum beat spectrum as a function of detection wavelength obtained from the PP data in Figure 3. The PP spectra shows two modes at 210 cm-1 and 575 cm-1 that are prominently seen on the $TT_{1}$ photoproduct and nearly absent in the ground electronic state. This can also be seen in Figure 5b which overlays the spectrally integrated GSB and ESA bands which are denoted by the gray vertical bands in Figure 5a. The ground state Raman spectrum for the NDI dimer, overlaid as left vertical panel, shows only weak Raman peaks corresponding to the quantum beats at $\sim$210 cm-1 and $\sim$575 cm-1 and consistent with their near absence in the GSB band observed in the PP experiments. Further, Figure S4a shows a similar plot as Figure 5a for the NDI monomer. Exactly similar trends are seen in the NDI monomer as well, that is, very weak to absent modes at $\sim$180 cm-1 and $\sim$580 cm-1 on the ground electronic state. The absence of these prominent modes on the ground electronic state of the NDI monomer is also consistent with the Raman spectrum of related NDI derivatives reported52 previously (see Figure 5 of ref. 52). Figure S4c also compares the mode frequencies between the NDI monomer and the dimer where shifts of 30 cm-1 for the 210 cm-1 mode and 5 cm-1 for the 575 cm-1 are seen. These shifts likely arise due to the constrained skeleton of the dimer.